OpenStudy (anonymous):

Which of the following could be solved using the percent equation: 14.4 = p% x 18?

OpenStudy (anonymous):

Answers: (A) Victor's breakfast cereal has fourteen point four mg of Iron. What percentage of the Recommended Daily Allowance (RDA) of Iron is in his cereal if the RDA is eighteen mg per day? (B) Victor's breakfast cereal has fourteen point four mg of Iron, which is eighteen percent of the Recommended Daily Allowance (RDA) of Iron. What is the RDA of Iron? (C) Victor's breakfast cereal has eighteen mg of Iron. What percentage of the Recommended Daily Allowance (RDA) of Iron is in his cereal if the RDA is fourteen point four mg per day? (D) Victor's breakfast cereal has eighteen mg of Iron, which is fourteen percent of the Recommended Daily Allowance (RDA) of Iron. What is the RDA of Iron?

OpenStudy (anonymous):

@e.cociuba

OpenStudy (e.cociuba):

Which of the following could be solved using the percent equation: 14.4 = 0.8 • b? 14.4=0.8•b 14.4=0.8b DIVISION is your answer. Divide by 0.8 on each side. b=18 A) Victor's breakfast cereal has 0.8 mg of Iron. What percentage of the Recommended Daily Allowance (RDA) of Iron is in his cereal if the RDA is 14.4 mg per day? Simply divide 0.8 for the amount of iron in Victor's breakfast cereal by 14.4 for the RDA of iron, then multiply that number by 100 since all percents are based on and out of 100. 0.8/14.4=0.055555555555556 0.055555555555555556*100=5.555555555555… of the recommended daily amount. B) Victor's breakfast cereal has 14.4 mg of Iron, which is 80% of the Recommended Daily Allowance (RDA) of Iron. What is the RDA of Iron? Simply multiply .80 times 14.4 to find the RDA of iron since all percents are based on and out of 100. .80*14.4=11.52 mg of iron for the Recommended daily amount C) Victor's breakfast cereal has 14.4 mg of Iron. What percentage of the Recommended Daily Allowance (RDA) of Iron is in his cereal if the RDA is 80 mg per day? Simply divide 14.4 for his breakfast cereal by 80 for the total cereal, then multiply that number by 100 since all percents are based on 100. 14.4/80=.18 .18*100=18% for the RDA of iron in his cereal if the RDA is 80 mg per day. D) Victor's breakfast cereal has 0.8 mg of Iron, which is 14.4% of the Recommended Daily Allowance (RDA) of Iron. What is the RDA of Iron? Simply divide 100 by 14.4, then multiply that number by 0.8 to find the RDA of iron. 100=14.4=6.944444444444444 6.9444444444444444444444*0.8=5.55555555… mg for the RDA of iron.